OpenStudy (anonymous):

The slope of a line is 11 and the point (14,19) is on the line. Express the line in slope- intercept form

OpenStudy (johnweldon1993):

Slope intercept form is \[\large y = mx + b\] where m = slope b = the y-intercept What you are given is a slope (m)...and a point...(x,y) so we are potentially solving for the y-intercept here plug in everything... \[\large 19 = 14(11) + b\] can you solve for 'b' ?

OpenStudy (anonymous):

the formula is y-y1+m(x-X1)

OpenStudy (johnweldon1993):

If you want to use that, thats fine too...that is point-slope form though... so if you want to set it up like that... \[\large y - 19 = (11)(x - 14)\] Now just rearrange and solve for 'y'

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

m=slope

OpenStudy (johnweldon1993):

Alright...so there were 2 ways to solve this...you could do the way I posted above...but you were also right in saying that you can use point slope form to solve this...either way will come out to the same answer

OpenStudy (johnweldon1993):

so lets use your point slope form equation \[\large y - y_1 = m(x - x_1)\]

OpenStudy (johnweldon1993):

You have a point....(14,19) this will be your (x1,y1) and you have a slope...this will be your 'm'

OpenStudy (anonymous):

i got y-19+11x-154

OpenStudy (anonymous):

m=slope so plug the slope into the m slot.

OpenStudy (johnweldon1993):

Right @Kyria just remember it should be y - 19 = 11x - 154 the minus sign does not go away \[\large y - 19 = 11(x - 14)\] becomes \[\large y - 19 = 11x - 154\] Now you want y by itself...so what do you do to that 19?

OpenStudy (anonymous):

y=11x+173

OpenStudy (anonymous):

Correct

OpenStudy (johnweldon1993):

Close...but remember it is -154 + 19...

OpenStudy (anonymous):

-135

OpenStudy (johnweldon1993):

Right...so it is \[\large y = 11x - 135\]

OpenStudy (johnweldon1993):

Great job @Kyria

OpenStudy (anonymous):

thanks i may need yhu help again

OpenStudy (anonymous):

stay in touch

OpenStudy (johnweldon1993):

No problem, If I'm not on your question feel free to tag me or message me :)

OpenStudy (anonymous):

thanks